OpenStudy (anonymous):

Solve: -2log8(x+1)=-8

OpenStudy (mertsj):

\[\log_{8}(x+1)=4 \]

OpenStudy (mertsj):

Now put that into exponential form.

OpenStudy (anonymous):

8^4= x+1?

OpenStudy (mertsj):

yep. Now solve it.

OpenStudy (anonymous):

x=4095

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

Thank you! Could you please help me with two more real quick?

OpenStudy (mertsj):

Possibly.